A brand of uncooked spaghetti comes in a box that is a rectangle prism with a lengthof 9 inches, width of 2 inches and a height of 1 1/2 inches. What is the surface area
1
Expert's answer
2020-02-20T09:56:34-0500

Solution. According to the conditions of the problem L=9 inches is length; W=2 inches is width; h=1.5 inches is height. For the rectangle prism the surface area is equal to


"A=2\times LW+2\times LH+2\times WH"

"A=2(9\times2+9\times 1.5+2\times 1.5)=69 in^2"

Answer.

"A=69 in^2"